# Language-Agnostic Modeling of Source Reliability on Wikipedia

### v2

@@ 5. Discussion
In this study, we have proposed and assessed editing behaviour-based features for modeling the reliability of web domains used as references on Wikipedia. Editing behaviour is not the only source of useful information for language-agnostic modeling of source reliability. Links between articles (already used for topical clustering of Wikipedia articles (Johnson et al., 2021)) may provide semantically meaningful context for references, whereas users’ reading sessions (used for entity linking (Gerlach et al., 2021)) may give controversiality signals from the perspective of the audience. Future work will also take a page from the social media literature and build editor-resource networks (Shu et al., 2019) that would capture structural peculiarities of topics, editors, and resources associated with low-quality content.
Apart from finding that the proposed features capture editorial behaviour within a dataset spanning one topic and language, we also observe that they retain useful information during the model adaptation across languages. Although English has been the lingua franca in NLP, having the best-developed tools and resources, we show that adaptation from the English-language dataset to other languages is largely inferior to using a model that combines information from many languages. This improvement is especially noticeable in performance on low-resource languages. These findings are aligned with limitations posed by Das et al. (2024) suggesting that, although the features are language-agnostic, limiting data selection to one language may limit the breadth of the captured behaviours and their applicability to new settings. As multilingual embedding models have gained popularity (Zhang et al., 2023a), it remains to be seen whether the latest developments in NLP will benefit the low-resource languages (Zhang et al., 2023b), and what the impact will be on Wikipedia specifically. [^9] https://en.wikipedia.org/wiki/Wikipedia:Why_you_shouldn%27t_write_articles_with_ChatGPT,_according_to_ChatGPT The need to monitor possible changes in the editorial standards and accompanying quality of new references calls for further development of features reflecting these behaviors, such as those described in the present study.
@@
The present study has multiple limitations. Using language-agnostic approaches may be seen as a limitation, as it does not take advantage of the predictive power of language-dependent features. However, despite the significant advances with multilingual NLP resources such as mBERT (Devlin et al., 2018), the motivation for our model was precisely to be ready for use in the 300+ language editions of Wikipedia, especially those that are less privileged and under-resourced.
Second, while the dataset contains thousands of articles and cumulatively millions of revisions, it spans only a fraction of Wikipedia. Despite choosing topical foci such that different reference quality and time scales, they may not capture the entire diversity of subject matter of this massive collaborative effort. For instance, differentiating topics by whether the topic is currently controversial may reveal shifts in editorial handling of sources.
Third, the selection of the English perennial sources list as the gold standard disadvantaged other languages. Unfortunately, only a few other languages have a similar resource and they are significantly less extensive and conclusive (Baigutanova et al., 2023b). Given the global role played historically by the English edition of Wikipedia (Hale, 2014), our selection provided the best choice to reach the maximum potential coverage across other language editions. In the future, we aim to use the scores computed by the adapted models as initial input for future collaboration with the editors of these language editions to create language-specific source quality guidelines.
@@
This study, and online content moderation in general, have important ethical aspects that should be kept in mind both by the platforms and the misinformation researchers.
First, this study has a direct impact on the access equity to encyclopedic resources in low-resource languages. Currently, the lack of resources such as the perennial source list for these versions of Wikipedia limits the extent to which our model is able to reflect the standards of that linguistic community. The domain scores provided by the models adapted from other (higher resource) languages that we propose here could be a starting point for bootstrapping the editorial discussion of the quality of sources used in low-resource languages.
Second, the process of determining the quality of sources bears scrutiny. In the news domain, the quality of information is often determined by professional fact-checkers. However, on Wikipedia, the editors from the community define the status of the sources via a deliberation process, guided by the community standards. Any potential biases in the judgment of these editors or failings in the deliberative process would be reflected in the quality of these labels (for instance, according to Kharazian et al. (2023), some Wikipedia editions were dominated for a time by a “small group of administrators who introduced far-right bias and outright disinformation”). Generally, in 2014 Shaw and Hill (2014) find that the Wikipedia contributors self-organize into oligarchic structures, though in 2021 Rijshouwer et al. (2023) find “strong counter-tendencies” to editorial power concentration. Our proposed method may reveal community behaviors which are not necessarily explicitly stated, but which shape the platform’s use of outside sources, potentially aiding the community to self-monitor the use of sources.

### v3

@@ 5. Discussion
In this study, we have proposed and assessed editing behavior-based features for modeling the reliability of web domains used as references on Wikipedia. We show that these models can be a part of a language-agnostic toolkit applicable both to high- and low-resource languages. As such, this work extends previous efforts to use the edit history to measure the controversiality of Wikipedia articles (Borra et al., 2015). Editing behavior is not the only source of useful information for language-agnostic modeling. Links between articles (already used for topical clustering of Wikipedia articles (Johnson et al., 2021)) may provide semantically meaningful context for references, whereas users’ reading sessions (used for entity linking (Gerlach et al., 2021)) may give controversiality signals from the perspective of the audience. Future work could also take a page from the social media literature and build editor-resource networks (Shu et al., 2019) to capture structural peculiarities of topics, editors, and resources associated with low-quality content. Additionally, automated feature discovery methods, such as self-supervised or representation learning on structured edit data, could help uncover latent behavioral patterns not easily captured by manually designed features. However, this would come at the cost of losing interpretability, a key value of our current approach.
Apart from finding that the proposed features capture editing behavior within a dataset spanning one topic and language, we also observe that they retain useful information during the model adaptation across languages. Although English has been the lingua franca in NLP, having the best-developed tools and resources, we show that adaptation from the English-language dataset to other languages is largely inferior to using a model that combines information from many languages. This improvement is especially noticeable in performance on low-resource languages, where native models often perform no better than random baselines, due to insufficient revision and citation activity for effective learning. However, when training on data pooled from all languages and applying quantile normalization to account for resource differences, the model significantly outperforms random baselines and matches or outperforms native models in most low-resource cases. This strategy highlights a practical and scalable pathway to empower under-resourced Wikipedia communities, even in the absence of extensive local data or tooling.
These findings are aligned with limitations identified by Das et al. (2024), who suggest that, although language-agnostic features are used, limiting data selection to one language may limit the breadth of the captured behaviors and their applicability to new settings. As multilingual embedding models have gained popularity (Zhang et al., 2023a), it remains to be seen to what extent the latest developments in NLP will benefit the low-resource languages (Zhang et al., 2023b), and what the impact will be on Wikipedia specifically (Johnson et al., 2024). In this context, the need to monitor possible changes in the editorial standards and citation practices across languages becomes especially important. Community efforts could prioritize on developing robust cross-lingual transfer strategies, rather than relying solely on local data that may be too sparse to train reliable predictors
@@
The present study has multiple limitations. First, using language-agnostic approaches may be seen as a limitation, as it does not take advantage of the predictive power of language-dependent features. However, despite the significant advances with multilingual NLP resources such as mBERT (Devlin et al., 2019), the motivation for our model was precisely to be ready for use in the 300+ language editions of Wikipedia, especially those that are less privileged and under-resourced.
Second, while the dataset contains thousands of articles and cumulatively millions of revisions, it spans only a fraction of Wikipedia. Despite choosing topical foci such that they have different reference quality and time scales, they may not capture the entire diversity of subject matter of this massive collaborative effort. For instance, differentiating topics by whether the topic is currently controversial may reveal shifts in editorial handling of sources. Additionally, future versions of our modeling approach could incorporate newly generated data by implementing re-training strategies using updated activity data, including changes in the classification of web domain reliability scores from the perennial source list.
Third, our approach may inherently favor generalist sources over topic-specific ones. Because several behavioral features (e.g., the number of articles or revisions in which a source domain appears) capture overall prevalence within Wikipedia, broadly used domains tend to accumulate stronger permanence and visibility signals. Conversely, specialized domains, despite being authoritative within their fields, may appear in fewer contexts and thus receive lower reliability estimates. This limitation suggests a potential bias toward sources covering a wider topical range. Its presence could be investigated in future research and mitigated by normalizing the corresponding features by topical diversity.
Fourth, the selection of the English perennial sources list as the gold standard disadvantaged other languages. Unfortunately, only a few other languages have a similar resource, and they are significantly less extensive and conclusive (Baigutanova et al., 2023b). Given the global role played historically by the English edition of Wikipedia (Hale, 2014), our selection provided the best choice to reach the maximum potential coverage across other language editions. However, the adaptation of the domain list from English may not coincide with possibly different cultural or local interpretations of quality by the editors from other language communities. In such cases, it may be preferable to apply NLP transfer methods (Das et al., 2025) that leverage higher-resource language editions to support lower-resource ones, particularly when they are comparable in the domain of source reliability. Our study, which employs language-agnostic features, can serve as a foundation for identifying appropriate language pairs for transfer in future work.
Fifth, our model relies on a fixed ground truth, consistent with prior work (e.g., (Baigutanova et al., 2023a, b)). However, perennial source lists evolve over time as editorial consensus changes and new sources are evaluated. While we treat it as static for this study, future work should incorporate time-aware labels and explore retraining strategies to adapt to such changes and maintain alignment with current community standards.
@@
This study, and online content moderation in general, have important ethical aspects that should be kept in mind by both the platforms and the misinformation researchers.
First, this study has a direct impact on the access equity to encyclopedic resources in low-resource languages. Currently, the lack of resources, such as the perennial source list for these versions of Wikipedia, limits the extent to which our model is able to reflect the standards of that linguistic community. The domain scores provided by the models adapted from other (higher resource) languages that we propose here could be a starting point for bootstrapping the editorial discussion of the quality of sources used in low-resource languages.
Second, the process of determining the quality of sources bears scrutiny. In the news domain, the quality of information is often determined by professional fact-checkers. However, on Wikipedia, the editors from the community define the status of the sources via a deliberation process, guided by the community standards. Any potential biases in the judgment of these editors or failings in the deliberative process would be reflected in the quality of these labels (for instance, according to Kharazian et al. (2024), some Wikipedia editions were dominated for a time by a “small group of administrators who introduced far-right bias and outright disinformation”). Generally, in 2014 Shaw and Hill (2014) concluded that the Wikipedia contributors self-organize into oligarchic structures, though in 2021 Rijshouwer et al. (2023) found “strong counter-tendencies” to editorial power concentration. Our proposed method may reveal community behaviors that are not necessarily explicitly stated, but which shape the platform’s use of outside sources, potentially aiding the community to self-monitor the use of sources.
@@
Acknowledgements.
This work is partially supported by a grant from the Credibility Coalition, a project of Meedan and Hacks/Hackers. Authors also acknowledge support from the Lagrange Project of the Institute for Scientific Interchange Foundation (ISI Foundation), funded by Fondazione Cassa di Risparmio di Torino (Fondazione CRT). Earlier versions of this paper benefited from valuable feedback from Diego Sáez Trumper, Francisco Navas, Aitolkyn Baigutanova, and attendees of Wiki Workshop 2024.
Appendix A Appendix A: Feature list
We provide here a more detailed list of the 52 features used by our models.
A.1. Popularity features
  1. (1)
    $N_{\mathrm{articles}}$: Number of articles a domain has appeared in (sensitive to the size of the dataset).
  2. (2)
    $\overline{N_{articles}}$: Number of articles a domain has appeared in, normalized by the number of articles in the dataset.
  3. (3)
    $CurrN_{articles}$: Number of articles a domain is used in at collection time.
  4. (4)
    $\overline{CurrN_{articles}}$: Number of articles a domain is used in at collection time, normalized by the total number of articles in the dataset.
A.2. Permanence features
  1. (5)
    $\Sigma Perm_{d}$: Permanence, how long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in days.
  2. (6)
    $\Sigma Perm_{r}$: Permanence, how long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured by his number of revisions.
  3. (7)
    $\Sigma CurrPerm_{d}$: How long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in days, but considering only the articles where the domain is currently used.
  4. (8)
    $\Sigma CurrPerm_{r}$: How long a domain has been used in an article (not necessarily consecutively), summed across all articles and measured in number of revisions, but considering only the articles where the domain is currently used.
  5. (9)
    $\overline{\Sigma Perm_{d}}$: Sum of all permanences for a domain in all articles, normalized by the sum of the ages of all articles, measured by the number of days.
  6. (10)
    $\overline{\Sigma Perm_{r}}$: Sum of all permanences for a domain in all articles, normalized by the sum of the ages of all articles, measured by the number of revisions.
  7. (11)
    $\overline{\Sigma CurrPerm_{d}}$: Sum of all permanences for a domain in articles where the domain was used at collection time, normalized by the sum of the ages of all articles, measured by the number of days.
  8. (12)
    $\overline{\Sigma CurrPerm_{r}}$: Sum of all permanences for a domain in articles where the domain was used at collection time, normalized by the sum of the ages of all articles, measured by the number of revisions.
  9. (13)
    $\langle Perm_{d}\rangle$: The average permanence of a domain (over all articles where it was used), measured by the number of days.
  10. (14)
    $\langle Perm_{r}\rangle$: The average permanence of a domain (over all articles where it was used), measured by the number of revisions.
  11. (15)
    $\langle SelfPerm_{d}\rangle$, : Self-permanence: permanence divided by the number of days since the first time the domains were added to the article, then averaged article-wise.
  12. (16)
    $\langle SelfPerm_{r}\rangle$: Self-permanence: permanence divided by the number of revisions since the first time the domains were added to the article, then averaged article-wise.
  13. (17)
    $\Sigma age_{d}$: The sum of ages of the domain over all the articles in a dataset, where the age of a domain in an article is the number of days since a URL from that domain has been first added to the article.
  14. (18)
    $\Sigma age_{r}$: The sum of ages of the domain over all the articles in a dataset, where the age of a domain in an article is the number of revisions since a URL from that domain has been first added to the article.
  15. (19)
    $\langle age_{d}\rangle$: The average age of a domain over all the articles it appears at least once, measured by the number of days.
  16. (20)
    $\langle age_{r}\rangle$: The average age of a domain over all the articles it appears at least once, measured by the number of revisions.
A.3. User-based features
  1. (21)
    $U_{add}$ Number of unique users that have added a domain.
  2. (22)
    $U_{start}$ Number of unique users that have added a domain for the first time on a page.
  3. (23)
    $U_{rem}$ Number of unique users that have removed a domain.
  4. (24)
    $U_{end}$ Number of unique users that have removed a domain for the last time on a page.
  5. (25)
    $R_{add}$ Number of unique registered users that have added a domain.
  6. (26)
    $R_{start}$ Number of unique registered users that have added a domain for the first time on a page.
  7. (27)
    $R_{rem}$ Number of unique registered users that have removed a domain.
  8. (28)
    $R_{end}$ Number of unique registered users that have removed a domain for the last time on a page.
  9. (29)
    $\overline{U_{add}}$ Number of unique users that have added a domain, divided by the total number of unique users in the dataset.
  10. (30)
    $\overline{U_{start}}$ Number of unique users that have added a domain for the first time on a page, divided by the total number of unique users in the dataset.
  11. (31)
    $\overline{U_{rem}}$: Number of unique users that have removed a domain, divided by the total number of unique users in the dataset.
  12. (32)
    $\overline{U_{end}}$: Number of unique users that have removed a domain for the last time on a page, divided by the total number of unique users in the dataset.
  13. (33)
    $\overline{R_{add}}$: Number of unique registered users that have added a domain, divided by the total number of unique users in the dataset.
  14. (34)
    $\overline{R_{start}}$: Number of unique registered users that have added a domain for the first time on a page, divided by the total number of unique users in the dataset.
  15. (35)
    $\overline{R_{rem}}$: Number of unique registered users that have removed a domain, divided by the total number of unique users in the dataset.
  16. (36)
    $\overline{R_{end}}$: Number of unique registered users that have removed a domain for the last time on a page, divided by the total number of unique users in the dataset.
  17. (37)
    $\langle{U_{add}}\rangle$: Average per page number of unique users that have added a domain.
  18. (38)
    $\langle{U_{start}}\rangle$: Average per page number of unique users that have added a domain for the first time on a page.
  19. (39)
    $\langle{U_{rem}}\rangle$: Average per page number of unique users that have removed a domain.
  20. (40)
    $\langle{U_{end}}\rangle$: Average per page number of unique users that have removed a domain for the last time on a page.
  21. (41)
    $\langle{R_{add}}\rangle$: Average per page number of unique registered users that have added a domain.
  22. (42)
    $\langle{R_{start}}\rangle$: Average per page number of unique registered users that have added a domain for the first time on a page.
  23. (43)
    $\langle{R_{rem}}\rangle$: Average per page number of unique registered users that have removed a domain
  24. (44)
    $\langle{R_{end}}\rangle$: Average per page number of unique registered users that have removed a domain for the last time on a page.
  25. (45)
    $Ratio(R_{add}/U_{add})$: The proportion of registered vs. all users that ever added a domain.
  26. (46)
    $Ratio(R_{start}/U_{start})$: The proportion of registered vs. all users that ever added a domain for the first time on a page.
  27. (47)
    $Ratio(R_{rem}/U_{rem})$: The proportion of registered vs. all users that ever removed a domain.
  28. (48)
    $Ratio(R_{end}/U_{end})$: The proportion of registered vs. all users that ever removed a domain for the last time on a page.
  29. (49)
    $Proba(R_{add})$: Probability that, when a domain is added, this is done by the registered users.
  30. (50)
    $Proba(R_{start})$: Probability that, when a domain is added for the first time on a page, this is done by the registered users.
  31. (51)
    $Proba(R_{rem})$: Probability that, when a domain is removed, this is done by the registered users.
  32. (52)
    $Proba(R_{end})$: Probability that, when a domain is removed for the last time on a page, this is done by the registered users.
Appendix B Appendix B: Medians of resource-wise statistics
Table A1. Statistics of the topic-language datasets used in the study grouped by high, mid, and low resource languages. All numbers are medians (apart from the number of languages). Perennial Domains (P. Domains) refers to the number of domains for the perennial source list in English found in the different datasets. The corresponding mean values can be found in Table 1.
Topic Langs Articles Revs Revs/Ar URLs URLs/Ar Domains P. Domains
High Climate change 7 1162 251085 62 34352 18 11992 191
COVID-19 7 595 136470 49 29826 21 6031 185
Biology 7 8228 388729 17 32315 3 8377 127
History 7 4839 417275 40 30957 5 9324 146
Media 7 3542 398138 46 43849 8 12441 240
Mid Climate change 36 325 23000 27 4399 9 1991 85
COVID-19 37 182 10026 19 4891 9 1326 99
Biology 37 2053 57234 11 5005 2 1643 59
History 37 1687 70988 26 5500 3 1901 62
Media 37 799 30098 20 5347 4 1991 123
Low Climate change 56 100 2403 12 633 6 373 25
COVID-19 75 24 518 9 466 9 214 31
Biology 41 440 $49$ 7 541 3 301 14
History 55 421 8020 13 449 3 256 16
Media 70 73 1063 6 263 3 148 22
Appendix C APPENDIX C: Language-specific dataset sizes
Table A2. Number of articles per language editions in the different datasets.
Language Climate change COVID-19 Biology History Media
de 479 274 463 549 472
en 975 856 931 1092 1460
es 339 258 397 461 509
fr 412 262 388 546 476
ja 217 145 230 246 382
pt 223 125 248 258 296
ru 257 162 317 435 377
ar 195 148 196 258 189
az 62 43 88 126 73
bg 109 44 141 175 116
bn 82 51 92 95 80
ca 176 94 217 242 148
cs 168 79 188 203 143
da 141 43 143 161 119
el 104 55 112 211 95
et 101 58 105 122 73
eu 104 36 94 144 68
fa 156 89 171 229 168
fi 184 68 184 177 178
he 184 146 230 301 225
hi 98 42 78 76 58
hr 99 33 104 148 83
hu 151 61 205 227 186
hy 56 39 70 112 83
id 136 83 129 154 122
it 259 161 289 426 427
ko 139 106 153 204 204
ms 73 63 82 91 67
nb 174 72 170 183 145
nl 238 114 270 295 233
pl 185 89 233 279 236
ro 106 45 122 171 109
simple 136 53 140 137 118
sk 105 26 118 137 91
sl 106 39 99 156 65
sr 115 59 129 230 126
sv 187 79 204 221 201
th 107 61 123 138 125
tr 157 89 160 247 182
uk 162 212 173 194 150
uz 36 39 33 84 42
vi 137 217 145 161 145
zh 205 246 212 250 326
$k$ 0 17 55 49 34
af 72 28 71 102 46
als 33 0 0 58 0
as 21 13 23 18 13
ast 84 20 88 130 84
ba 28 6 0 36 14
bcl 21 9 16 27 12
be 56 27 61 90 48
be-x-old 45 14 0 77 42
bs 74 21 71 118 52
ckb 31 16 0 44 38
cv 25 0 0 0 25
cy 63 23 556 161 55
dag 29 0 0 0 0
eo 104 26 121 137 79
ga 43 24 44 76 47
gl 108 34 127 152 81
gu 29 6 0 0 9
ha 32 23 0 20 8
ia 31 11 0 0 0
ig 17 0 7 10 9
jv 46 15 63 74 32
ka 66 24 76 113 77
kn 51 28 40 33 35
ku 30 13 0 58 31
la 69 19 89 130 54
lb 33 0 0 0 53
lt 102 26 120 129 79
lv 80 27 95 93 63
mk 72 18 73 124 59
ml 70 39 83 81 49
mn 43 0 42 0 18
mni 13 0 0 0 0
mnw 4 5 0 0 0
mr 63 20 0 0 33
my 37 18 0 29 16
ne 35 16 42 0 23
$n$ 87 14 83 101 59
oc 55 0 0 98 38
or 16 18 0 0 14
pa 34 24 29 36 25
pnb 28 9 38 38 16
ps 21 9 0 23 0
rm 6 0 0 0 0
rw 14 0 19 0 0
sh 67 20 62 100 52
si 51 16 32 24 16
sq 56 36 50 104 60
su 30 19 44 0 13
sw 44 14 47 74 31
ta 90 33 88 68 57
te 49 33 61 40 47
tg 30 9 0 0 23
tl 47 40 65 76 41
ur 57 37 58 99 42
war 34 0 0 48 17
zh-yue 58 54 62 59 49
am 0 6 0 0 0
ary 0 4 0 0 0
arz 0 23 32 79 38
awa 0 2 0 0 0
bjn 0 11 0 0 0
bo 0 4 0 0 0
br 0 13 0 90 41
bug 0 2 0 0 0
bxr 0 7 0 0 0
ceb 0 7 0 0 0
crh 0 3 0 0 0
dty 0 2 0 0 0
gn 0 6 0 0 0
gpe 0 1 0 0 0
ht 0 26 0 0 0
hyw 0 3 0 33 18
km 0 8 0 9 0
mg 0 7 0 0 0
min 0 17 0 0 0
pap 0 12 0 0 0
sco 0 13 31 31 24
shn 0 5 0 0 0
so 0 9 0 0 0
tk 0 6 0 0 0
ts 0 5 0 0 0
vo 0 3 0 0 12
wuu 0 15 0 0 0
xmf 0 10 0 0 17
yo 0 24 0 0 0
zh-min-nan 0 14 0 39 25
is 0 0 75 91 50
ky 0 0 21 20 12
pam 0 0 27 0 0
$t$ 0 0 26 57 14
azb 0 0 0 32 12
fo 0 0 0 26 16
fy 0 0 0 81 32
io 0 0 0 79 0
tum 0 0 0 2 0
an 0 0 0 0 35
frp 0 0 0 0 6
glk 0 0 0 0 3
lo 0 0 0 0 5
mai 0 0 0 0 8
nds 0 0 0 0 28
qu 0 0 0 0 15
sat 0 0 0 0 12
vec 0 0 0 0 19
Table A3. Number of revisions per language editions in the different datasets.
Language Climate change COVID-19 Biology History Media
de 344149 195448 388729 549521 398138
en 1516501 1036268 2392272 2213972 3850137
es 251085 142608 412920 417275 447902
fr 261336 136470 401974 570207 405992
ja 59235 37293 97684 117733 222585
pt 82137 32740 178778 153532 154532
ru 90492 56512 230846 348758 257507
ar 64147 52221 123245 164152 73088
az 5396 3275 37635 34400 6289
bg 14688 4606 50747 77809 24812
bn 9138 6567 12759 20227 7907
ca 43861 13878 219014 168055 45265
cs 37535 15008 70480 94004 36957
da 21730 4267 33759 54452 21325
el 15873 7175 16066 93809 12233
et 12000 6702 21939 28705 6906
eu 12238 2086 82623 45030 6865
fa 34394 14882 65428 89813 59550
fi 41846 10081 76296 70988 58117
he 45760 36839 84560 153974 74220
hi 10574 3708 7683 12406 4471
hr 11084 2842 21316 40023 9546
hu 27013 14401 93431 100103 50952
hy 5149 4177 10735 29745 12491
id 23373 14547 111305 55091 32762
it 102487 46926 204053 375945 347881
ko 23966 28724 57234 84491 88325
ms 6906 7698 11679 20204 7208
nb 36680 10026 81352 78627 40317
nl 78355 25913 303604 171942 93987
pl 45284 22540 160392 173581 116880
ro 14046 7347 25254 60565 19300
simple 30367 8317 31239 33583 23031
sk 11945 2913 18615 40322 9525
sl 12320 2265 12780 39772 5602
sr 14728 6756 46083 102224 20365
sv 49949 11079 264477 114709 71292
th 12084 11217 22152 36162 30098
tr 34878 19488 61404 121987 56558
uk 33040 55024 105302 102155 45483
uz 2067 3116 2961 17800 2682
vi 22628 66920 407504 45861 30883
zh 53587 131339 131500 122649 182721
$k$ 0 517 6564 5187 1523
af 6384 2827 12553 20980 3448
als 1320 0 0 4863 0
as 814 466 1025 590 270
ast 7812 997 40421 36452 8628
ba 1316 50 0 2004 449
bcl 1036 182 439 1082 158
be 3892 1473 10090 23364 2899
be-x-old 2385 518 0 12592 1722
bs 5029 1324 7127 30720 2982
ckb 1397 564 0 3609 2026
cv 860 0 0 0 746
cy 4720 1280 521785 50767 9683
dag 790 0 0 0 0
eo 11720 1920 29807 42486 7768
ga 2570 1830 4293 8020 2404
gl 13814 2608 32768 44564 11127
gu 1191 311 0 0 157
ha 1676 638 0 628 131
ia 1437 265 0 0 0
ig 884 0 116 284 166
jv 2325 492 $49$ 12716 1271
ka 5100 1312 11997 33511 8342
kn 2643 1366 2085 1482 1593
ku 1659 656 0 4968 1020
la 5306 749 11783 40788 3739
lb 1925 0 0 0 2999
lt 11336 $16$ 32199 34848 7025
lv 7180 4234 13517 12402 5800
mk 6096 811 8133 35526 4200
ml 6461 2373 14983 7912 3545
mn 2421 0 2011 0 474
mni 385 0 0 0 0
mnw 160 217 0 0 0
mr 4801 1140 0 0 1494
my 1469 1358 0 1317 430
ne 1383 487 3090 0 684
$n$ 8312 1801 10620 20234 4729
oc 4226 0 0 19691 1618
or 313 916 0 0 286
pa 1609 825 1498 2143 832
pnb 1076 100 4884 4351 290
ps 693 144 0 850 0
rm 148 0 0 0 0
rw 390 0 400 0 0
sh 4436 619 13554 28448 5352
si 2728 421 1249 869 469
sq 3919 2758 3169 24390 4022
su 1328 523 1885 0 389
sw 3057 416 2273 13290 1027
ta 10484 2321 12039 6974 4530
te 3599 3142 4649 2462 3206
tg $14$ 331 0 0 637
tl 3262 2685 6835 11473 2215
ur 4132 2055 3941 18574 2528
war 1436 0 0 7500 419
zh-yue 3844 6256 6081 7762 3438
am 0 115 0 0 0
ary 0 539 0 0 0
arz 0 1221 12611 18006 10561
awa 0 48 0 0 0
bjn 0 153 0 0 0
bo 0 57 0 0 0
br 0 530 0 22230 2022
bug 0 7 0 0 0
bxr 0 238 0 0 0
ceb 0 125 0 0 0
crh 0 50 0 0 0
dty 0 16 0 0 0
gn 0 89 0 0 0
gpe 0 31 0 0 0
ht 0 1304 0 0 0
hyw 0 118 0 1401 401
km 0 464 0 134 0
mg 0 138 0 0 0
min 0 518 0 0 0
pap 0 387 0 0 0
sco 0 223 1636 1609 689
shn 0 143 0 0 0
so 0 164 0 0 0
tk 0 95 0 0 0
ts 0 98 0 0 0
vo 0 33 0 0 412
wuu 0 341 0 0 0
xmf 0 370 0 0 342
yo 0 1353 0 0 0
zh-min-nan 0 393 0 3607 889
is 0 0 7778 12301 2541
ky 0 0 1317 705 297
pam 0 0 798 0 0
$t$ 0 0 2568 11215 459
azb 0 0 0 3732 1496
fo 0 0 0 1252 307
fy 0 0 0 12889 1099
io 0 0 0 15974 0
tum 0 0 0 16 0
an 0 0 0 0 1325
frp 0 0 0 0 36
glk 0 0 0 0 61
lo 0 0 0 0 47
mai 0 0 0 0 104
nds 0 0 0 0 801
qu 0 0 0 0 435
sat 0 0 0 0 133
vec 0 0 0 0 569